OpenStudy (anonymous):

Find the logarithm log28^1

OpenStudy (australopithecus):

anything to the power of one is the same thing so you have log(28)

OpenStudy (lgbasallote):

\(\log 28^1???\)

OpenStudy (australopithecus):

type it into your calculator or wolfram alpha

OpenStudy (australopithecus):

28^(1) = 28 x^(1) = x 32423942^(1) = 32423942 2^(2) = (2)(2) x^(2) = x(x) etc.. Hope this helps